In the sequence 123, 133, 113, 123, what is the next number?
Answer by MathLover1(20850)


given:
the sequence 123, 133, 113, 123
look at the numbers in your sequence
the difference between first one and second one is +10...123+10=133
the difference between second one and third one is -20...133-20=113
the difference between third one and fourth one is +10...113+10=123
following this pattern, next number in this sequence will be 123 -20=103